OpenStudy (anonymous):

Two cars leave an intersection at the same time. One heads due south and the other heads due east. Later, the two cars are exactly 50 mi apart. The car headed south has traveled 10 mi farther than the car going east. How far has the car going east traveled? Please Help!!!!!!!!!

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Let x be the distance that the car moving East has traveled. Then x+10 is the distance that the car moving South has traveled. Solve the following for x:\[x^2+(x+10)^2=50^2 \]x = 30